# Coprocessor Actor Critic: A Model-Based Reinforcement Learning Approach For Adaptive Brain Stimulation

### v1

@@ 1 Introduction
A neural coprocessor is a form of brain-computer interface (BCI) that can transmit signals to and from the brain (Rao, 2019; Oehrn et al., 2023). These interfaces can be used to treat a variety of neurological conditions by influencing abnormal neural activity (Lozano et al., 2019; Little et al., 2013b). In patients who suffer from conditions such as Parkinson’s disease and dystonia, brain stimulation has the ability to steer neural activity towards activity regions which manifest in reduced disease symptoms (Hu & Stead, 2014; Groiss et al., 2009). Adaptive brain stimulation can be employed not just to guide neural activity towards specific activity patterns but can also aid impaired patients in accomplishing external task objectives (Bryan et al., 2022; Cunha et al., 2015). Stroke patients are one patient population that can benefit from this aspect of brain stimulation. Stroke patients suffer injury to the brain that often results in loss of motor control and an inability to complete basic tasks, such as reaching for and grasping an object (Hatem et al., 2016). Stroke patients often struggle with these seemingly simple motor tasks due to stroke-induced lesions in the brain that can interrupt the propagation of neurological signals within and between cortical modules (Ingwersen et al., 2021; Choi et al., 2023). Due to the resultant motor deficits, a patient may recognize the target location and intend to move their arm to the perceived position, but struggle to do so (Choi et al., 2023). Adaptive brain stimulation exhibits potential for reducing motor impairment and restoring lost function via adaptive stimulation, enabling these patients to operate more effectively in the world (Elias et al., 2018; Ganguly et al., 2022).
@@
In this work, we investigate the development of neural coprocessors to deliver adaptive brain stimulation for rehabilitation using an in silico model of brain injury. Neural coprocessors rely on artificial intelligence techniques to learn a brain stimulation policy that appropriately shapes neural activity based upon the current state of the patient (Rao, 2019). An effective coprocessor policy can compensate for lost mobility and paretic motor deficits post-stroke (Bryan et al., 2022). However, there are several challenges in developing an effective coprocessor policy. Because of the heterogeneity of patients’ brains, their disease manifestation, and the location of the stimuli, the optimal coprocessor policy is unique for each patient (Visanji et al., 2022). Furthermore, due to the complexity of the brain, closed-loop coprocessor policies are difficult to hand-engineer (Oehrn et al., 2023). Instead, individualized coprocessor policies may be learned through interaction with the patient to ensure that the coprocessor aligns seamlessly with the unique characteristics of the patient’s brain and condition. Due to the necessity for online learning, adaptive brain stimulation poses a compelling paradigm which can benefit from the application of reinforcement learning (RL) techniques.
@@
To learn how to act optimally in the world, we rely on a biomechanical simulator. Various physics based simulators (e.g., Caggiano et al. (2022) and Delp et al. (2007)) have been developed that enable the physiologically realistic simulation of human biomechanics. These simulations are capable of modeling both complex human physiology and dynamic environment interactions, such as the complexities encountered in our example reaching task. We leverage a biomechanical simulator to learn the value of a human executing a world action (e.g., moving their arm) in a given state (e.g., arm joint positions), for a given task objective (e.g., reaching a target location). That is, we employ a simulator to learn the Q-function for world actions.
@@ 2 Related Work
An alternative to MFRL is Model-Based Reinforcement Learning (MBRL). MBRL reduces training time and improves learning efficiency by utilizing a predictive dynamics model to learn an effective policy (Valencia et al., 2023; Janner et al., 2021). MBRL can typically be broken down into two steps: 1) dynamic model learning followed by 2) integration and planning (Moerland et al., 2022). The downside of MBRL is that the model used for planning may be inaccurate, thus producing suboptimal plans (Abbeel et al., 2006). Our approach minimizes this risk by leveraging a simulation-derived Q-function to guide model learning and sample stimulations that produce high-value environment actions, enabling us to robustly and efficiently learn the brain dynamics model.
@@ 4.2 Training Brain Model
We update the brain model via the mean-squared error loss between the predicted world action, $\hat{a}$ and the ground truth world action $a$. To effectively capture the complexity of the relationship between stimulations and world actions, we adopt a structure for $\widehat{F}_{\text{brain}}^{\,\phi}(s,\bar{a})$ akin to the model presented (Bryan et al., 2022). In Bryan et al. (2022), the authors leverage a neural network to learn the mapping from stimulations to world actions from a monkey stroke dataset and show that this model is able to effectively capture the effects of stimulation on the brain
@@ 5 Experiments
We compare CopAC to the popular MFRL approach, Soft Actor-Critic (SAC) (Haarnoja et al., 2018), which combines actor and critic networks with an entropy regularization term, promoting exploration in a stable and efficient manner. We choose to compare to SAC because actor-critic algorithms have been employed in prior work in learning a policy for closed-loop brain stimulation (Gao et al., 2023; Lu et al., 2020a). We additionally baseline against the MBRL approach Model-Based Policy Optimization (Janner et al., 2021), which trains a model of the environment and uses both real experience and simulated experience from the model to update its policy. To assess the effectiveness of our sampling policy and the importance of updating the world-action value model, we conduct an ablation experiment and compare against CopAC with a random sampling policy (instead of maximizing the Q-function) as well as CopAC without updating the world-action value model.
@@ 5.1 In Silico Evaluation Environments
Figure 2: This figure shows the brain stimulation domain for the MyoSim Arm Reach task. We model the biomechanics of the reaching tasks using the MyoSuite physics simulator (Caggiano et al., 2022). The brain of a stroke patient is modeled via the approach described by Michaels et al. (2020) and consists of the anterior intraparietal area (AIP), ventral premotor cortex (F5), and primary motor cortex (M1) modules. The coprocessor applies stimulation to the motor cortex (M1) which modifies the world action of the patient.
@@
Physiologically and Neurologically Realistic Stroke Domain: Drawing on prior work in neurophysiological and biomechanical modeling (Michaels et al., 2020; Caggiano et al., 2022), we introduce a novel simulation domain for evaluating adaptive brain stimulation policies in stroke patients (Figure 2). Such an in silico evaluation requires both a neurologically realistic human brain model of a stroke patient and a high-fidelity biomechanical simulator. To simulate the biomechanics of the human musculoskeletal system, we rely on MyoSuite (Caggiano et al., 2022), a cutting-edge simulator for biomechanical control problems based on the MuJoCo physics engine. The action space consists of individual muscle activations and the observation space consists of joint angles.
@@ Acknowledgements
This work was supported in part by a grant from the Weill Neurohub and by CoCoSys, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.

### v2

@@ 1 Introduction
A neural coprocessor is a form of brain-computer interface (BCI) that can transmit signals to and from the brain (Rao, 2019; Oehrn et al., 2023). These interfaces can be used to treat a variety of neurological conditions by influencing abnormal neural activity (Lozano et al., 2019; Little et al., 2013b). In patients who suffer from conditions such as Parkinson’s disease and dystonia, brain stimulation has the ability to steer neural activity towards activity regions which manifest in reduced disease symptoms (Hu & Stead, 2014; Groiss et al., 2009). Adaptive brain stimulation can be employed not just to guide neural activity towards specific activity patterns but can also aid impaired patients in accomplishing external task objectives (Bryan et al., 2023; Cunha et al., 2015). Stroke patients are one patient population that can benefit from this aspect of brain stimulation. Stroke patients suffer injury to the brain that often results in loss of motor control and an inability to complete basic tasks, such as reaching for and grasping an object (Hatem et al., 2016). Stroke patients often struggle with these seemingly simple motor tasks due to stroke-induced lesions in the brain that can interrupt the propagation of neurological signals within and between cortical modules (Ingwersen et al., 2021; Choi et al., 2023). Due to the resultant motor deficits, a patient may recognize the target location and intend to move their arm to the perceived position, but struggle to do so (Choi et al., 2023). Adaptive brain stimulation exhibits potential for reducing motor impairment and restoring lost function via adaptive stimulation, enabling these patients to operate more effectively in the world (Elias et al., 2018; Ganguly et al., 2022).
@@
In this work, we investigate the development of neural coprocessors to deliver adaptive brain stimulation for rehabilitation using an in silico model of brain injury. Neural coprocessors rely on artificial intelligence techniques to learn a brain stimulation policy that appropriately shapes neural activity based upon the current state of the patient (Rao, 2019). An effective coprocessor policy can compensate for lost mobility and paretic motor deficits post-stroke (Bryan et al., 2023). However, there are several challenges in developing an effective coprocessor policy. Because of the heterogeneity of patients’ brains, their disease manifestation, and the location of the stimuli, the optimal coprocessor policy is unique for each patient (Visanji et al., 2022). Furthermore, due to the complexity of the brain, closed-loop coprocessor policies are difficult to hand-engineer (Oehrn et al., 2023). Instead, individualized coprocessor policies may be learned through interaction with the patient to ensure that the coprocessor aligns seamlessly with the unique characteristics of the patient’s brain and condition. Due to the necessity for online learning, adaptive brain stimulation poses a compelling paradigm which can benefit from the application of reinforcement learning (RL) techniques.
@@
To learn how to act optimally in the world, we rely on a biomechanical simulator. Various physics based simulators (e.g., Wang et al. (2022) and Delp et al. (2007)) have been developed that enable the physiologically realistic simulation of human biomechanics. These simulations are capable of modeling both complex human physiology and dynamic environment interactions, such as the complexities encountered in our example reaching task. We leverage a biomechanical simulator to learn the value of a human executing a world action (e.g., moving their arm) in a given state (e.g., arm joint positions), for a given task objective (e.g., reaching a target location). That is, we employ a simulator to learn the Q-function for world actions.
@@ 2 Related Work
An alternative to MFRL is Model-Based Reinforcement Learning (MBRL). MBRL reduces training time and improves learning efficiency by utilizing a predictive dynamics model to learn an effective policy (Valencia et al., 2023; Janner et al., 2019). MBRL can typically be broken down into two steps: 1) dynamic model learning followed by 2) integration and planning (Moerland et al., 2023). The downside of MBRL is that the model used for planning may be inaccurate, thus producing suboptimal plans (Abbeel et al., 2006). Our approach minimizes this risk by leveraging a simulation-derived Q-function to guide model learning and sample stimulations that produce high-value environment actions, enabling us to robustly and efficiently learn the brain dynamics model.
@@ 4.2 Training Brain Model
We update the brain model via the mean-squared error loss between the predicted world action, $\hat{a}$ and the ground truth world action $a$. To effectively capture the complexity of the relationship between stimulations and world actions, we adopt a structure for $\widehat{F}_{\text{brain}}^{\,\phi}(s,\bar{a})$ akin to the model presented (Bryan et al., 2023). In Bryan et al. (2023), the authors leverage a neural network to learn the mapping from stimulations to world actions from a monkey stroke dataset and show that this model is able to effectively capture the effects of stimulation on the brain
@@ 5 Experiments
We compare CopAC to the popular MFRL approach, Soft Actor-Critic (SAC) (Haarnoja et al., 2018), which combines actor and critic networks with an entropy regularization term, promoting exploration in a stable and efficient manner. We choose to compare to SAC because actor-critic algorithms have been employed in prior work in learning a policy for closed-loop brain stimulation (Gao et al., 2023; Lu et al., 2020a). We additionally baseline against the MBRL approach Model-Based Policy Optimization (Janner et al., 2019), which trains a model of the environment and uses both real experience and simulated experience from the model to update its policy. To assess the effectiveness of our sampling policy and the importance of updating the world-action value model, we conduct an ablation experiment and compare against CopAC with a random sampling policy (instead of maximizing the Q-function) as well as CopAC without updating the world-action value model.
@@ 5.1 In Silico Evaluation Environments
Figure 2: This figure shows the brain stimulation domain for the MyoSim Arm Reach task. We model the biomechanics of the reaching tasks using the MyoSuite physics simulator (Wang et al., 2022). The brain of a stroke patient is modeled via the approach described by Michaels et al. (2020) and consists of the anterior intraparietal area (AIP), ventral premotor cortex (F5), and primary motor cortex (M1) modules. The coprocessor applies stimulation to the motor cortex (M1) which modifies the world action of the patient.
@@
Physiologically and Neurologically Realistic Stroke Domain: Drawing on prior work in neurophysiological and biomechanical modeling (Michaels et al., 2020; Wang et al., 2022), we introduce a novel simulation domain for evaluating adaptive brain stimulation policies in stroke patients (Figure 2). Such an in silico evaluation requires both a neurologically realistic human brain model of a stroke patient and a high-fidelity biomechanical simulator. To simulate the biomechanics of the human musculoskeletal system, we rely on MyoSuite (Wang et al., 2022), a cutting-edge simulator for biomechanical control problems based on the MuJoCo physics engine. The action space consists of individual muscle activations and the observation space consists of joint angles.
@@ Acknowledgements
This work was supported in part by a grant from NSF HCC, the Weill Neurohub, and by CoCoSys, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.